OpenStudy (anonymous):

which expression is equivalent to this expression? (5x^2-4x+5)-(3x^2-6x+2) A. 2x^2+2x+3 B. 15x^2-10x+7 C. 2x^2-2x+3 D. 2x^2-10x+7

OpenStudy (anonymous):

The answer would be choice number A

OpenStudy (anonymous):

Reason being is because 5x^2 - 3x^2 = 2x^2 -4x-(-6x) = 2, because double negative makes a positive And then 5-2=3

OpenStudy (anonymous):

Ohh okay thank you

OpenStudy (anonymous):

No problem :)